OpenStudy (anonymous):

The domain of the function f(x) = 3x^3 is {2, 5}. What is the functions range? The range of the function f(x) = 3x^3 when its domain is {2, 5} is ... ?

OpenStudy (solomonzelman):

you basically just need to plug in 2 for x into the f(x), to find what f(2) is, and do the same thing to 5. plug in 5 into f(x), to find f(5).

OpenStudy (solomonzelman):

write these two values { f(2), f(5) } as a range corresponding to the domain of {2,5} but DON'T write just f(2) and f(5), put the actual (numerical) values of f(2) and f(5) into the range.

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x) = 3x^3 }\) if I were to find f(3), I would do this: \(\large\color{black}{ f3) = 3(3)^3=3\times 27=81 }\) So, \(\large\color{black}{ f(3) = 81 }\)

OpenStudy (solomonzelman):

Can you find the f(2), and f(5) the same way?

OpenStudy (anonymous):

Im sorry im still confused

OpenStudy (solomonzelman):

okay. can you find f(2)?

OpenStudy (anonymous):

how do i find it , I put f(2)=3x^3 .. how do I solve it from there ? .I dont have a calculator so its just me & my paper

OpenStudy (solomonzelman):

\(\large\color{black}{ f(2) =3(2)^3 }\)

OpenStudy (solomonzelman):

I am plugging in 2 instead of x, every where.... can you calculate the f(2) now?

OpenStudy (anonymous):

ohh , so ill be 6^3 Now ?

OpenStudy (solomonzelman):

you first do the exponent, and then multiply it times 3.

OpenStudy (solomonzelman):

\(\large\color{black}{ (2)^3 }\) is equal to ? then multiply THAT times 3. this will give you the value of \(\large\color{black}{3(2)^3 }\) (which is the f(2) ).

OpenStudy (anonymous):

when i multiply the exponent by 2 does it give me the exponent or I keep it the same nd change 2 ?

OpenStudy (solomonzelman):

what is 2^3, do you know?

OpenStudy (anonymous):

Its 24 right ? @SolomonZelman

OpenStudy (solomonzelman):

yes f(2) is 24.

OpenStudy (solomonzelman):

How about f(5)?

OpenStudy (solomonzelman):

plug in 5 for x, into f(x).

OpenStudy (solomonzelman):

just like we did with a 2 to find f(2).

OpenStudy (anonymous):

Okay I got 375

OpenStudy (anonymous):

so 3(2)^3 is 24 & 3(5)^3 is 375 . so whats next ? @SolomonZelman

OpenStudy (solomonzelman):

so you have your f(2)=24 and f(5)=375 you domain for f(x)=3x^3 was {2,5} and therefore the range is {f(2), f(5) } but instead of f(2) and f(5) put the values that we found for f(2) and f(5).

OpenStudy (anonymous):

oh okayy , Thankss for helping me

OpenStudy (solomonzelman):

anytime:)